If 3 times Jane's age, in years, is equal to 8 times Beth's age, in years, and the difference between their ages is 15 years. How old are Jane and Beth?
Answer by matthew_sessoms(39)

Jane = j
Beth = b


3j = 8b
j - b = 15


We have here a "system of equations." Lets solve for "j" in the second equation and then plug it in into the first equation...


j = 15 + b


3(15 + b) = 8b
45 + 3b = 8b
45 = 5b
9 = b


Now, plug "b" into either of the original equations. I chosed the second equation because it looks easier.


j - 9 = 15
j = 24


Jane's age = 24yrs.
Beth's age = 9yrs.

MS